# RepLoRA: Reparameterizing Low-Rank Adaptation via the Perspective of Mixture of Experts

### v1

@@ 1 Introduction
With the rapid growth in data availability and computational resources, large-scale models trained on extensive datasets have demonstrated remarkable generalization capabilities, enabling successful applications across language, vision, and multi-modal tasks [3, 41, 48]. However, fully fine-tuning such models for specific downstream tasks can be prohibitively expensive. To address this challenge, several parameter-efficient fine-tuning (PEFT) methods [11, 24, 16] have emerged, facilitating effective adaptation of large pre-trained models by adjusting a minimal set of parameters while keeping most of the backbone frozen. Among these methods, Low-Rank Adaptation (LoRA) [12] stands out for its simplicity and effectiveness and has been successfully applied across diverse domains [26, 39, 47, 30]. Despite its successes, theoretical understanding of LoRA has remained limited, hindering our ability to optimize its performance further.
Building on the recent finding [20] about the connection between attention mechanism [50] and the mixture of experts (MoE) [15, 17] models, we present a rigorous theoretical study demonstrating how LoRA can be interpreted within this new framework. Leveraging this perspective, we show that a straightforward reparameterization technique [29, 21], which represents low-rank matrices as the output of an MLP, can theoretically enhance the performance of LoRA. Specifically, our analysis reveals that this reparameterization can reduce the data needed to achieve a desired estimation error from an exponential scale to a polynomial scale, thereby substantially improving sample efficiency. Based on these insights, we introduce Reparameterized Low-Rank Adaptation (RepLoRA). This novel PEFT method reparameterizes low-rank matrices through a lightweight MLP.
@@
Organizations. The paper is organized as follows: Section 2 provides the background on LoRA and MoE. Section 3 establishes the connection between LoRA and MoE. Section 4 presents our theoretical analysis, including the statistical benefits of reparameterizing LoRA. Building on these insights, Section 5 presents our method, RepLoRA. To demonstrate the effectiveness of RepLoRA, Section 6 presents the experimental results. Finally, Section 7 concludes the paper.
@@
Notation. We denote $[n]:=\{1,2,\ldots,n\}$ for any $n\geq 1$. Furthermore, the notation $|S|$ denotes the cardinality of a set $S$. Given a vector $u:=(u_{1},u_{2},\ldots,u_{d})\in\mathbb{R}^{d}$ and a coefficient $\alpha:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, we have the following notations $u^{\alpha}:=u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}\ldots u_{d}^{\alpha_{d}}$, $|u|:=u_{1}+u_{2}+\ldots+u_{d}$ and $\alpha!:=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$. Additionally, $\|u\|$ stands for the Euclidean norm of vector $u$. For positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates $a_{n}/b_{n}$ is stochastically bounded.
@@ 2 Preliminaries
Multi-head Self-attention. We begin by revisiting the architecture of the multi-head self-attention (MSA) layer in Transformer [50, 3]. Let ${\bm{X}}=[{\bm{x}}_{1},\dots,{\bm{x}}_{N}]^{\top}\in\mathbb{R}^{N\times d}$ denote an input sequence of embeddings, where $N$ is the sequence length and $d$ denotes the embedding dimension. The MSA layer processes this sequence as follows:
@@
where we define each attention head by ${\bm{h}}_{i}=\mathrm{Attention}({\bm{X}}\bm{W}_{i}^{Q},{\bm{X}}\bm{W}_{i}^{K},{\bm{X}}\bm{W}_{i}^{V})$ for $i\in[m]$. Here, ${\bm{X}}_{Q}=({\bm{X}}\bm{W}_{1}^{Q},\ldots,{\bm{X}}\bm{W}_{m}^{Q})$, $\mathbb{X}_{K}=({\bm{X}}\bm{W}_{1}^{K},\ldots,{\bm{X}}\bm{W}_{m}^{K})$, and $\mathbb{X}_{V}=({\bm{X}}\bm{W}_{1}^{V},\ldots,{\bm{X}}\bm{W}_{m}^{V})$ are the query, key, and value matrices, respectively. Furthermore, $m$ is the number of heads, and $\bm{W}^{O}\in\mathbb{R}^{md_{v}\times d}$ is the output projection matrix. Each attention head ${\bm{h}}_{i}$ is parameterized by $\bm{W}_{i}^{Q}\in\mathbb{R}^{d\times d_{k}},\bm{W}_{i}^{K}\in\mathbb{R}^{d\times d_{k}},\text{ and }\bm{W}_{i}^{V}\in\mathbb{R}^{d\times d_{v}}$, with $d_{k}=d_{v}=\frac{d}{m}$.
Low-Rank Adaptation. LoRA [12] has emerged as an efficient method for adapting large pre-trained transformer models to downstream tasks. Building upon the hypothesis that the updates during fine-tuning exhibit a low “intrinsic rank”, LoRA proposes to fine-tune the transformer architectures’ linear layers by incrementally updating the pre-trained weights with the product of two low-rank matrices. For a given pre-trained weight matrix $\bm{W}_{0}\in\mathbb{R}^{m\times n}$, LoRA represents its update as $\Delta\bm{W}={\bm{B}}{\bm{A}}$, where ${\bm{B}}\in\mathbb{R}^{m\times r}$, and ${\bm{A}}\in\mathbb{R}^{r\times n}$ with $r\ll\min\{m,n\}$. Consequently, the output of the fine-tuned model is:
@@
During training, $\bm{W}_{0}$ remains fixed, while ${\bm{A}}$ and ${\bm{B}}$ are updated. Typically, LoRA is used to adjust the linear layers that generate transformer models’ queries and values (or keys, queries, and values). In line with prior work [12, 31, 53], here we fine-tune the query and value projection matrices, leading to the following output expression:
@@
Mixture of Experts. A mixture of experts (MoE) [15, 17] model consists of $N$ expert networks, $f_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d_{v}}$ for $i\in[N]$, and a gating function $G:\mathbb{R}^{d}\to\mathbb{R}^{N}$ that allocates contributions of each expert based on the input ${\bm{x}}$ to the model. The output of the MoE model is given by:
@@
where $G({\bm{x}})=\mathrm{softmax}(s_{1}({\bm{x}}),\cdots,s_{N}({\bm{x}}))$, and $s_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is a score function. In the subsequent sections, we discuss how MoE relates to LoRA and provide a theoretical analysis of our method.
@@ 3 LoRA from the perspective of MoE
Prior work [20, 21] has shown that each attention head in the MSA layer can be viewed as a specialized architecture of multiple MoE models. Specifically, from Eq. (1), consider the output of the $l$-th head ${\bm{h}}_{l}=[{\bm{h}}_{l,1},\dots,{\bm{h}}_{l,N}]^{\top}\in\mathbb{R}^{N\times d_{v}}$. Let $\mathbb{X}=\left[{\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{N}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ denote the concatenated input embeddings. We then define $N$ experts $f_{j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$ encoded within the MSA layer as:
@@
for $i\in[N]$ and $j\in[N]$. Then, we can formulate each output vector ${\bm{h}}_{l,i}$ as the output of an MoE model, using the experts and score functions above:
@@
Within a pre-trained MSA layer, all parameters of these experts and score functions $\bm{W}_{l}^{Q}$, $\bm{W}_{l}^{K}$, and $\bm{W}_{l}^{V}$ remain fixed. When LoRA is applied, it refines these experts and score functions with updates:
@@ 4 Theoretical Analysis of LoRA: With and Without Reparameterization
Above, we assume that $\mathbb{X}_{1},\mathbb{X}_{2},\ldots,\mathbb{X}_{n}$ are i.i.d. samples from some probability distribution $\mu$ with bounded support. Meanwhile, the noise variables $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are independent and following Gaussian distributions such that $\mathbb{E}[{\varepsilon_{i}}|\mathbb{X}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|\mathbb{X}_{i})=\nu^{2}I_{\bar{d}}$ for all $i\in[n]$. Next, the regression function $f_{G_{*}}(\cdot)$ takes the form of an MoE model with $L$ unknown experts, that is,
$f_{G_{*}}(\mathbb{X}):=\sum_{j=1}^{L}\frac{\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{B}}_{Q,j}^{*}{\bm{A}}_{Q,j}^{*}){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{j})}{D_{f}(\mathbb{X})}\cdot({\bm{M}}^{0}_{V}+{\bm{B}}_{V,j}^{*}{\bm{A}}_{V,j}^{*})\mathbb{X},$ (11)
where we denote the following normalization term
@@
With versus Without Reparametrization. Subsequently, we establish the convergence rates of estimating the unknown low-rank matrices $\{{\bm{B}}_{Q,j^{\prime}}^{*},{\bm{A}}_{Q,j^{\prime}}^{*},{\bm{B}}_{V,j^{\prime}}^{*},{\bm{A}}_{V,j^{\prime}}^{*}\}_{j^{\prime}=1}^{L}$ under two scenarios, namely without shared structures among these low-rank matrices (equivalently, without reparametrization) in Section 4.1 and with shared structures among these low-rank matrices (equivalently, with reparametrization) in Section 4.2. Our ultimate goal is to demonstrate that the sample efficiency of estimating these low-rank matrices under the shared structures setting is much better than that under the non-shared structures setting with a given error $\epsilon>0$. The theory sheds light on our Reparameterized LoRA (RepLoRA) design in Section 5.
@@ 4.1 Without Reparametrization: Suboptimal Sample Complexity
We begin our convergence analysis for low-rank matrices with the scenario where the LoRA reparametrization is absent. It is worth noting that we can estimate those unknown matrices via estimating the ground-truth mixing measure $G_{*}$; for that sake, we utilize the least square method [49] to obtain the following estimator:
$\widehat{G}_{n}:=\textnormal{arg}\min_{G\in\mathcal{G}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{\|}{\bm{Y}}_{i}-f_{G}(\mathbb{X}_{i})\Big{\|}^{2},$ (12)
where the set of all mixing measures with at most $L^{\prime}$ atoms is defined as
$\mathcal{G}_{L^{\prime}}(\Theta):=\{G=\sum_{j^{\prime}=1}^{\ell}\exp(c_{j^{\prime}})\delta_{({\bm{B}}_{Q,j^{\prime}},{\bm{A}}_{Q,j^{\prime}},{\bm{B}}_{V,j^{\prime}},{\bm{A}}_{V,j^{\prime}})}:1\leq\ell\leq L^{\prime},\ (c_{j^{\prime}},{\bm{B}}_{Q,j^{\prime}},{\bm{A}}_{Q,j^{\prime}},{\bm{B}}_{V,j^{\prime}},{\bm{A}}_{V,j^{\prime}})\in\Theta\}.$
As the number of ground-truth experts $L$ is typically unknown in practice, we assume that the number of fitted experts $L^{\prime}$ is large enough for $L^{\prime}>L$. Then, to determine the convergence rates of the estimator $\widehat{G}_{n}$, we use a loss function built upon the concept of Voronoi cells [33].
@@
$\mathcal{V}_{j}:=\{i\in[L^{\prime}]:\|{\bm{H}}_{i}-{\bm{H}}^{*}_{j}\|\leq\|{\bm{H}}_{i}-{\bm{H}}^{*}_{\ell}\|,\forall\ell\neq j\},$
@@
$\mathcal{D}_{1,r}(G,G_{*}):=\sum_{j^{\prime}=1}^{L}\Big{|}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})-\exp(c^{*}_{j^{\prime}})\Big{|}+\sum_{j^{\prime}=1}^{L}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})(\|\Delta{\bm{B}}_{Q,ij^{\prime}}\|^{r}+\|\Delta{\bm{A}}_{Q,ij^{\prime}}\|^{r}$
@@
for $r\in\mathbb{N}$, where $\Delta{\bm{H}}_{ij^{\prime}}:={\bm{H}}_{i}-{\bm{H}}^{*}_{j^{\prime}}$ for any $i,j^{\prime}$. Now, we are ready to study the sample efficiency of LoRA without reparametrization in Theorem 4.1, whose proof is deferred to Appendix A.1.
@@
For any $r\in\mathbb{N}$, the following bound holds:

### v2

@@ Abstract
Machine Learning, ICML, LoRA, Mixture of Experts, Reparameterize
@@ 1 Introduction
With the rapid growth in data availability and computational resources, large-scale models trained on extensive datasets have demonstrated remarkable generalization capabilities, enabling successful applications across language, vision, and multi-modal tasks (Dosovitskiy, 2021; Radford et al., 2021; Touvron et al., 2023). However, fully fine-tuning such models for specific downstream tasks can be prohibitively expensive. To address this challenge, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Jia et al., 2022) have emerged, facilitating effective adaptation of large pre-trained models by adjusting a minimal set of parameters while keeping most of the backbone frozen. Among these methods, Low-Rank Adaptation (LoRA) (Hu et al., 2022) stands out for its simplicity and effectiveness and has been successfully applied across diverse domains (Li et al., 2022; Qin et al., 2023; Taori et al., 2023; Liu et al., 2024a). Despite its successes, theoretical understanding of LoRA has remained limited, hindering our ability to optimize its performance further.
Building on the recent finding (Le et al., 2024) about the connection between attention mechanism (Vaswani, 2017) and the mixture of experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1994) models, we present a rigorous theoretical study demonstrating how LoRA can be interpreted within this new framework. Leveraging this perspective, we show that a straightforward reparameterization technique (Li & Liang, 2021; Le et al., 2025), which represents low-rank matrices as the output of an MLP, can theoretically enhance the performance of LoRA. Specifically, our analysis reveals that this reparameterization can reduce the data needed to achieve a desired estimation error from an exponential scale to a polynomial scale, thereby substantially improving sample efficiency. Based on these insights, we introduce Reparameterized Low-Rank Adaptation (RepLoRA) - a novel PEFT method reparameterizes low-rank matrices through a lightweight MLP.
@@
Organization. The paper is organized as follows: Section 3 provides the background on LoRA and MoE. Section 4 establishes the connection between LoRA and MoE. Section 5 presents our theoretical analysis, including the statistical benefits of reparameterizing LoRA. Building on these insights, Section 6 presents our method, RepLoRA. To demonstrate the effectiveness of RepLoRA, Section 7 presents the experimental results. Finally, Section 8 concludes the paper.
@@
Notation. For any $n\in\mathbb{N}$, let $[n]:=\{1,2,\ldots,n\}$. For any set $S$, $|S|$ denotes its cardinality. Given a vector $u:=(u_{1},u_{2},\ldots,u_{d})\in\mathbb{R}^{d}$ and $\alpha:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, we define $u^{\alpha}=u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}\ldots u_{d}^{\alpha_{d}}$, $|u|:=u_{1}+u_{2}+\ldots+u_{d}$ and $\alpha!:=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$, while $\|u\|$ stands for its Euclidean norm. For positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates $a_{n}/b_{n}$ is stochastically bounded.
2 Related Works
Parameter-Efficient Fine-tuning (PEFT).
With the recent rise of large models, PEFT methods are growing in popularity for their ability to fine-tune large-scale models by training a relatively small number of parameters for adapting to specific downstream tasks. Existing PEFT methods can be divided into three categories. The first category is referred to as adapter-based methods, which introduce additional trainable parameters to the frozen backbone. For example, Series Adapter (Houlsby et al., 2019) proposes adding linear modules in sequence to the existing layer, or Parallel Adapter (He et al., 2022) proposes integrating these modules in parallel. The second category of PEFT methods is Prompt-based methods that add extra trainable soft tokens, referred to as prompts, to the input (Lester et al., 2021; Razdaibiedina et al., 2023; Wang et al., 2023). A weakness of these methods is that they increase inference latency compared to the original model.
Low-Rank Adaptation (Hu et al., 2022). LoRA and its variants are among the third category of the PEFT method, which is well-known for its simplicity and for not adding extra inference burden. To fine-tune the linear layers of a large model, LoRA applies low-rank matrices to approximate the weight changes and then merges them with the pre-trained weights for inference. A recent variant of LoRA is DoRA (Liu et al., 2024b), which proposes to decompose the weight change into a learnable magnitude and directional component. Another example is AdaLoRA (Zhang et al., 2023), which parameterizes the incremental updates in the form of singular value decomposition and prunes less significant singular values for more efficient updates. Orthogonal Fine-tuning (OFT) (Qiu et al., 2023) exploits the orthogonal factorization to fine-tune diffusion models. Recently, VeRA (Kopiczko et al., 2024) significantly reduces the number of trainable parameters compared to LoRA by using learnable scaling vectors to adjust a shared pair of frozen random matrices across layers. Notably, our method, which will be proposed in the following few sections, also falls within this category, and we validate its efficacy alongside LoRA and its variants through theoretical analysis and comprehensive experimentation.
Mixture of Experts. Expanding upon the foundational principles of mixture models (Jacobs et al., 1991; Jordan & Jacobs, 1994), earlier research (Eigen et al., 2014; Shazeer et al., 2017) established the Mixture of Experts (MoE) layer as a crucial mechanism for effectively scaling model capacity. Over time, MoE models have gained significant recognition due to their versatility across multiple domains, including large language models (Du et al., 2022; Zhou et al., 2023), computer vision (Riquelme et al., 2021; Puigcerver et al., 2024), multimodal learning (Han et al., 2024), and multi-task learning (Ma et al., 2018). Recent studies have focused on analyzing the convergence rates of expert estimation in MoE models, exploring different assumptions and configurations related to gating mechanisms and expert functions. For instance, in the context of softmax gating, Nguyen et al. (2023, 2024a, 2024b) revealed that the expert estimation rates are shaped by the solvability of polynomial systems arising from interactions between gating and expert parameters. More recently, Nguyen et al. (2024d, c, 2025) employed least squares estimation to establish an identifiability condition for expert functions, particularly in feedforward networks with nonlinear activation functions. These findings indicate that under these conditions, estimation rates improve significantly compared to models relying on polynomial experts.
@@ 3 Preliminaries
Multi-head Self-attention. We begin by revisiting the architecture of the multi-head self-attention (MSA) layer in Transformer (Vaswani, 2017; Dosovitskiy, 2021). Let ${\bm{X}}=[{\bm{x}}_{1},\dots,{\bm{x}}_{N}]^{\top}\in\mathbb{R}^{N\times d}$ denote an input sequence of embeddings, where $N$ is the sequence length and $d$ denotes the embedding dimension. The MSA layer processes this sequence as follows:
@@
where each attention head is defined by ${\bm{h}}_{i}=\mathrm{Attention}({\bm{X}}\bm{W}_{i}^{Q},{\bm{X}}\bm{W}_{i}^{K},{\bm{X}}\bm{W}_{i}^{V})$ for $i\in[m]$. Here, ${\bm{X}}_{Q}=({\bm{X}}\bm{W}_{1}^{Q},\ldots,{\bm{X}}\bm{W}_{m}^{Q})$, $\mathbb{X}_{K}=({\bm{X}}\bm{W}_{1}^{K},\ldots,{\bm{X}}\bm{W}_{m}^{K})$, and $\mathbb{X}_{V}=({\bm{X}}\bm{W}_{1}^{V},\ldots,{\bm{X}}\bm{W}_{m}^{V})$ are the query, key, and value matrices, respectively. Furthermore, $m$ is the number of heads, and $\bm{W}^{O}\in\mathbb{R}^{md_{v}\times d}$ is the output projection matrix. Each attention head ${\bm{h}}_{i}$ is parameterized by $\bm{W}_{i}^{Q}\in\mathbb{R}^{d\times d_{k}},\bm{W}_{i}^{K}\in\mathbb{R}^{d\times d_{k}},\text{ and }\bm{W}_{i}^{V}\in\mathbb{R}^{d\times d_{v}}$, with $d_{k}=d_{v}=\frac{d}{m}$.
Low-Rank Adaptation. LoRA (Hu et al., 2022) has emerged as an efficient method for adapting large pre-trained transformer models to downstream tasks. Building upon the hypothesis that the updates during fine-tuning exhibit a low “intrinsic rank”, LoRA proposes to fine-tune the transformer architectures’ linear layers by incrementally updating the pre-trained weights with the product of two low-rank matrices. For a given pre-trained weight matrix $\bm{W}_{0}\in\mathbb{R}^{m\times n}$, LoRA represents its update as $\Delta\bm{W}={\bm{B}}{\bm{A}}$, where ${\bm{B}}\in\mathbb{R}^{m\times r}$, and ${\bm{A}}\in\mathbb{R}^{r\times n}$ with $r\ll\min\{m,n\}$. Consequently, the output of the fine-tuned model is:
@@
During training, $\bm{W}_{0}$ remains fixed, while ${\bm{A}}$ and ${\bm{B}}$ are updated. Typically, LoRA adjusts the linear layers that generate transformer models’ queries and values (or keys, queries, and values). In line with prior work (Hu et al., 2022; Liu et al., 2024b; Xin et al., 2024), here we fine-tune the query and value projection matrices, leading to the following output expression:
@@
Mixture of Experts. A mixture of experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1994) model consists of $N$ expert networks, $f_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d_{v}}$ for $i\in[N]$, and a gating function $G:\mathbb{R}^{d}\to\mathbb{R}^{N}$ that allocates contributions of each expert based on the input ${\bm{x}}$ to the model. The output of the MoE model is given by:
@@
where $G({\bm{x}})=\mathrm{softmax}(s_{1}({\bm{x}}),\cdots,s_{N}({\bm{x}}))$, and $s_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is a score function. In the subsequent sections, we discuss how MoE relates to LoRA and provide a theoretical analysis of the proposed method.
@@ 4 LoRA from the perspective of MoE
Prior work (Le et al., 2024, 2025) has shown that each attention head in the MSA layer can be viewed as a specialized architecture of multiple MoE models. Specifically, from Eq. (1), consider the output of the $l$-th head ${\bm{h}}_{l}=[{\bm{h}}_{l,1},\dots,{\bm{h}}_{l,N}]^{\top}\in\mathbb{R}^{N\times d_{v}}$. Let $\mathbb{X}=\left[{\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{N}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ denote the concatenated input embeddings. We then define $N$ experts $f_{j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$ encoded within the MSA layer as follows:
@@
for $i\in[N]$ and $j\in[N]$. Consequently, each output vector ${\bm{h}}_{l,i}$ can be formulated as the result of an MoE model, utilizing the experts and score functions defined above:
@@
Within a pre-trained MSA layer, all parameters of these experts and score functions $\bm{W}_{l}^{Q}$, $\bm{W}_{l}^{K}$, and $\bm{W}_{l}^{V}$ remain fixed. When LoRA is applied, it refines these experts and score functions with low-rank updates:
@@ 5 Theoretical Analysis of LoRA: With and Without Reparameterization
Above, we assume that $\mathbb{X}_{1},\mathbb{X}_{2},\ldots,\mathbb{X}_{n}$ are i.i.d. samples from some probability distribution $\mu$ with bounded support. Meanwhile, $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|\mathbb{X}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|\mathbb{X}_{i})=\nu^{2}I_{\bar{d}}$ for all $i\in[n]$. Next, the regression function $f_{G_{*}}(\cdot)$ takes the form of an MoE model with $L$ unknown experts, that is,
$f_{G_{*}}(\mathbb{X}):=\sum_{j=1}^{L}\frac{\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{B}}_{Q,j}^{*}{\bm{A}}_{Q,j}^{*}){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{j})}{D_{f}(\mathbb{X})}$
$\hskip 85.35826pt\cdot({\bm{M}}^{0}_{V}+{\bm{B}}_{V,j}^{*}{\bm{A}}_{V,j}^{*})\mathbb{X},$ (11)
where we denote
@@
With versus Without Reparametrization. Subsequently, we establish the convergence rates of estimating the unknown low-rank matrices $\{{\bm{B}}_{Q,j^{\prime}}^{*},{\bm{A}}_{Q,j^{\prime}}^{*},{\bm{B}}_{V,j^{\prime}}^{*},{\bm{A}}_{V,j^{\prime}}^{*}\}_{j^{\prime}=1}^{L}$ under two scenarios, namely without shared structures among these low-rank matrices (equivalently, without reparametrization) in Section 5.1 and with shared structures among these low-rank matrices (equivalently, with reparametrization) in Section 5.2. Our ultimate goal is to demonstrate that sample efficiency of estimating these low-rank matrices under the shared structures setting is much better than that under the non-shared structures setting with a given error $\epsilon>0$. The theory sheds light on our design of Reparameterized LoRA (RepLoRA) in Section 6.
@@ 5.1 Without Reparametrization: Suboptimal Sample Complexity
We begin our convergence analysis for low-rank matrices with the scenario where the LoRA reparametrization is absent. It is worth noting that we can estimate those unknown matrices via estimating the ground-truth mixing measure $G_{*}$, For that sake, we utilize the least square method (van de Geer, 2000) to obtain the following estimator:
$\widehat{G}_{n}\in\textnormal{arg}\min_{G\in\mathcal{G}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{(}{\bm{Y}}_{i}-f_{G}(\mathbb{X}_{i})\Big{)}^{2},$ (12)
where we denote by $\mathcal{G}_{L^{\prime}}(\Theta):=\{G=\sum_{j^{\prime}=1}^{\ell}\exp(c_{j^{\prime}})\delta_{({\bm{B}}_{Q,j^{\prime}},{\bm{A}}_{Q,j^{\prime}},{\bm{B}}_{V,j^{\prime}},{\bm{A}}_{V,j^{\prime}})}:1\leq\ell\leq L^{\prime},\ (c_{j^{\prime}},{\bm{B}}_{Q,j^{\prime}},{\bm{A}}_{Q,j^{\prime}},{\bm{B}}_{V,j^{\prime}},{\bm{A}}_{V,j^{\prime}})\in\Theta\}$ the set of all mixing measures with at most $L^{\prime}$ atoms. As the number of ground-truth experts $L$ is typically unknown in practice, we assume that the number of fitted experts $L^{\prime}$ is large enough such that $L^{\prime}>L$. Then, to determine the convergence rates of the estimator $\widehat{G}_{n}$, we use a loss function built upon the concept of Voronoi cells (Manole & Ho, 2022).
@@
$\mathcal{V}_{j}:=\{i\in[L^{\prime}]:\|{\bm{Z}}_{i}-{\bm{Z}}^{*}_{j}\|\leq\|{\bm{Z}}_{i}-{\bm{Z}}^{*}_{\ell}\|,\forall\ell\neq j\},$
@@
$\mathcal{D}_{1,r}(G,G_{*}):=\sum_{j^{\prime}=1}^{L}\Big{|}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})-\exp(c^{*}_{j^{\prime}})\Big{|}$
$+\sum_{j^{\prime}=1}^{L}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})(\|\Delta{\bm{B}}_{Q,ij^{\prime}}\|^{r}+\|\Delta{\bm{A}}_{Q,ij^{\prime}}\|^{r}$
@@
for $r\in\mathbb{N}$, where $\Delta{\bm{H}}_{ij^{\prime}}:={\bm{H}}_{i}-{\bm{H}}^{*}_{j^{\prime}}$ for any $i,j^{\prime}$. Now, we are ready to study the sample efficiency of LoRA without reparametrization in Theorem 5.1 whose proof is deferred to Appendix A.1.
@@
The following bound holds for any $r\in\mathbb{N}$: